# Dynamic Pricing with Adversarially-Censored Demands

### v1

@@ 4 Algorithm
In this section, we present our core algorithm,C20CB, which stands for a Closest-To-Zero Confidence Bound strategy that proposes asymptotically optimal prices over differentiated inventory levels and censoring effects. The structure of C20CB is displayed as Algorithm 1.
Algorithm 1 C20CB:Closest-To-Zero Confidence Bound
1:  Input: Quantities $\tau$ and $a_{\max},b_{\min},b_{\max},p_{\max},c$, coefficients $\gamma_{0},C_{a},C_{b},C_{F},C_{G},C_{N},C_{\tau}$.
2:  STAGE 1: Pure Exploration
3:  for  $t=1,2,\ldots,\tau$ do
4:     Sample and propose a price $p_{t}\sim U[0,p_{\max}]$ uniformly at random.
5:     Observe demand $D_{t}$ and indicators $e_{i,t}:=\mathds{1}[D_{t}\geq\frac{i\cdot\gamma_{t}+(4-i)\gamma_{0}}{4}],i=1,2,3$.
6:  end for
7:  Estimate
$\hat{b}=$ $\frac{1}{4p_{\max}\cdot\frac{1}{\tau}\sum_{t=1}^{\tau}\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}}$ (4)
$\hat{a}=$ $\frac{1}{\tau}\sum_{t=1}^{\tau}(\hat{b}p_{\max}e_{3,t}+\frac{3\gamma_{t}+\gamma_{0}}{4}).$
8:  STAGE 2: Optimistic Acting
9:  Denote $\Delta:=(C_{a}+C_{b}\cdot p_{\max})\cdot\frac{1}{\sqrt{\tau}}$.
@@
10:  Define $M:=\lfloor\frac{c}{2\Delta}\rfloor$ and $w_{k}:=2k\Delta,k=-M,-M+1,\ldots,-1,0,1,\ldots,M$.
11:  for $t=1,2,\ldots,2M+1$ do
12:     Let $k_{t}:=-M-1+t$ and propose $p_{t}=\frac{w_{k_{t}}-(\gamma_{t}-\hat{a})}{\hat{b}}$.
13:     Observe $D_{t}$ and $\mathds{1}_{t}:=\mathds{1}[D_{t}<\gamma_{t}]$.
14:     Initialize $F_{k_{t}}\leftarrow\mathds{1}_{t},N_{k_{t}}\leftarrow 1,G_{k_{t}}\leftarrow D_{t}-\gamma_{t}+c,\Delta_{k_{t}}\leftarrow C_{F}\cdot b_{\max}p_{\max}+C_{G}+C_{b}\cdot\frac{1}{\sqrt{\tau}}$.
15:  end for
16:  for $t=1,2,\ldots,T-\tau-(2M+1)$ do
17:     if $\gamma_{t}\geq\frac{\hat{a}+C_{a}\cdot\frac{1}{\sqrt{\tau}}}{2}+c$ then
18:        Propose $p_{t}=\frac{\hat{a}}{2\hat{b}}$ and continue to $t+1$ (without recording feedback).
19:     end if
20:     Initialize $k_{t}\leftarrow M,\rho_{t}\leftarrow+\infty$.
21:     for $k=M,M-1,\ldots,-M+1,-M$ do
22:        Denote $p_{k,t}:=\frac{w_{k}-(\gamma_{t}-\hat{a})}{\hat{b}}$ and $\hat{r}_{k,t}:=\gamma_{0}-c+G_{k}-\hat{b}\cdot p_{k,t}\cdot F_{k}$.
23:        if $\hat{r}_{k,t}-\Delta_{k}\leq 0\leq\hat{r}_{k,t}+\Delta_{k}$ then
24:           Update $k_{t}\leftarrow k,\rho_{t}\leftarrow 0$, and Break.
25:        end if
26:        Let $\rho_{k,t}:=\min\{|\hat{r}_{k,t}-\Delta_{k}|,|\hat{r}_{k,t}+\Delta_{k}|\}$.
27:        if $\rho_{k,t}<\rho_{t}$ then
28:           Update $\rho_{t}\leftarrow\rho_{k,t}$ and $k_{t}\leftarrow k$.
29:        end if
30:     end for
31:     if $\hat{r}_{k,t}-\Delta_{k}>0,\forall k=-M,-M+1,\ldots,M-1,M$ then
32:        Propose $p_{t}=\frac{\hat{a}}{2\hat{b}}.$
33:     else
34:        Propose $p_{t}=p_{k_{t},t}$.
35:     end if
36:     Observe $D_{t}$ and $\mathds{1}_{t}=\mathds{1}[D_{t}<\gamma_{t}]$.
37:     Update
@@ 4.1 Algorithm Design Overview
    Stage 1: Exploration: During the first $\tau=\sqrt{T}$ rounds, the seller (we) proposes uniformly random prices in the range of $[0,p_{\max}]$. By the end of Stage 1, we obtain $\hat{a}$ and $\hat{b}$ as plug-in estimators of $a$ and $b$ in the following stage.
@@
    Stage 2: Optimistic Decision: We estimate the derivatives of the revenue function at discretized prices $\{p_{k,t}\}$’s. For each $p_{k,t}$, we not only estimate $r^{\prime}_{t}(p_{k,t}$ but also maintain an error bar of that estimate. At each time $t$, we propose the price whose corresponding error bar covers 00 or closest to 00 if no covering exists.
Algorithm 1 exhibits several advantageous properties. It is suitable for processing streaming data as the constructions of $\hat{a},\hat{b},\hat{r}^{\prime}_{t}(\cdot)$ are updated incrementally with each new observation (including $e_{i,t},D_{t},\mathds{1}_{t}$) without the need of revisiting any historical data. Additionally, it consumes $\tilde{O}(T^{\frac{5}{4}})$ time complexity and $O(T^{\frac{1}{4}})$ extra space, which are plausible for large $T$. A potential risk of computation might arise on the calculation of $\hat{b}$, where $\sum_{t=1}^{\tau}e_{1,t}-e_{2,t}$ can be 0 with a small but nonzero probability. Although this event does not undermine the high-probability regret guarantee, it might still be harmful to the computational system for numerical experiments. To mitigate this incident in practice, we may either extend Stage 1 until one non-zero $e_{1,t}-e_{2,t}=1$ is observed, or restart Stage 1 at $t=\tau$.
In the following sections, we will introduce each stage and technical component in details.
@@ 4.2 Pure-Exploration to Estimate Parameters from Biased Observations
We incorporate a uniform-exploration phase for estimating $a$ and $b$ in our algorithm, bypassing the obstacle brought by demand censoring. This approach is supported by the following insight: When $Y$ is a uniformly distributed random variable within a closed interval $[L,R]$, and $X$ is another random variable, independent to $Y$ and also distributed within $[L,R]$, we have:
@@
Here the second step uses the Law of Total Expectation. Eq. 6 indicates that we can derive an unbiased estimator of $\mathbb{E}[X]$ through $\mathds{1}[Y\geq X]$ even in the absence of any direct observation of $X$. Looking back to our algorithm, when $p_{t}\sim U[0,p_{\max}]$, we have
$\mathbb{E}[e_{i,t}]$ $=\mathbb{E}[\mathds{1}[a-bp_{t}+N_{t}\geq\gamma_{i}]]=\mathbb{E}[\mathbb{E}[N_{t}\geq\gamma_{i}-a+bp_{t}|N_{t}]]=\mathbb{E}[\frac{N_{t}-\gamma_{i}+a}{bp_{\max}}]=\frac{a-\gamma_{i}}{bp_{\max}}.$ (7)
Here $\gamma_{i}:=\frac{i\gamma_{t}+(4-i)\gamma_{0}}{4}$. The last equality comes from $\mathbb{E}[N_{t}]=0$. By deploying different $\gamma_{i}$ at $i=1,2,3$, we can estimate $a$ and $b$ through the observations of $e_{i,t}$, effectively circumventing the censoring effect. A similar technique has been utilized by Fan et al., (2024) to construct an unbiased estimator of the valuations instead of the demands as we concern. However, their application of uniform exploration might be sub-optimal as they adopt an exploration-then-exploitation design in each epoch. On the contrary, our algorithm uses this uniform exploration merely as a trigger of further learning. Our tight regret bound indicates that uniform exploration can still contribute to an optimal algorithm for a broad range of online learning instances.
4.3 Optimistic Strategy to Balance Derivatives Estimates v.s. Revenue Loss
Figure 1: The price which C20CB proposes based on confidence bounds of $\hat{r}_{k,t}$: (a) If there exist prices whose error bar contain 00, then we propose the largest price among them. (b) If no error bar contains 00 but there does exist at least one below 00, we propose the price whose corresponding error bar is closest to 00. (c) If all error bars are above 00, then we propose $\frac{\hat{a}}{2\hat{b}}$ for pure exploitation.
We proceed to Stage 2 with $\hat{a}$ and $\hat{b}$ established, which indicates we have an estimate on the expected potential linear demand $d_{t}(p)=a-bp$. However, we are still unaware of the noise distribution, which is crucial for the current optimal price as the inventory level $\gamma_{t}$ partially censors the noise.
In order to balance the learning of noise distribution versus the loss of proposing sub-optimal prices, we adopt an optimistic which involves the following components:
@@
    We discretize the $[-c,c]$ domain of noise CDF $F(\cdot)$ and its intergration $G(\cdot)$ into small intervals of length $2\Delta$, with $\Delta=O(\frac{1}{T^{1/4}})$. At the center of each interval $k$ (which is $2k\Delta$) we maintain independent estimates of $F$ and $G$, including their expectations and high-confidence error bars.
@@
    At each time $t$, we construct a set of discrete prices $\{p_{k,t}\}_{k=-M}^{M}$ such that the quantity $\gamma_{t}-\hat{a}+\hat{b}p_{k,t}$ matches the center of the interval $k$ established above. According to Eq. 13, we construct an estimate of each $r^{\prime}_{t}(p_{k,t})$ with plug-in estimators $\hat{a},\hat{b}$ and discrete estimators of $F$ and $G$ for the specific Interval $k$. Again, the estimate includes its expectation and error bar.
@@
    Since the optimal price $p_{t}^{*}$ satisfies $r^{\prime}_{t}(p_{t}^{*})=0$, we identify the discrete price $p_{k,t}$ where the derivative estimate is "possibly 00" or "closest to 00". We illustrate this process in Figure 1, which includes the following three cases:
@@
      If there exists some $p_{k,t}$ such that the corresponding error bar (of its derivative estimate) contains 00, we propose the largest price satisfying this condition.
@@
      If there does not exist $p_{k,t}$ whose corresponding error bar contains 00 (but there does exist error bar below 00), we propose the price whose error bound is closest to 00.
@@
      If all error bars are above 00, indicating that the reward function is monotonically increasing over the “censoring area”, we propose $p_{t}=\frac{\hat{a}}{2\hat{b}}$ to exploit the non-censoring optimal price $\frac{a}{2b}$. In this case, we do not need to record any observations nor to update any parameter/estimate.
@@
Here we provide an intuition of the optimality: On the one hand, the width of interval can tolerate the error of mapping from $2k\Delta$ to $p_{k,t}$, and the Lipschitzness of $F$ and $G$ ensures that our estimate within each small interval is roughly correct. On the other hand, we can show that the closest-to-zero derivative estimate implies a closest-to-$p_{t}^{*}$ price according to some locally strong convexity in the neighborhood of $p_{t}^{*}$. As we have smoothness on the regret function, we suffer a quadratic loss at $(T\cdot\frac{1}{T^{1/4}})^{2}=O(\sqrt{T})$ cumulatively, which balances the loss of Stage 1 that costs $O(\tau)=O(\sqrt{T})$ as well. For a rigorous regret analysis, we kindly refer the readers to Section 5.
4.4 Technical Novelty.
To the best of our knowledge, we are the first to introduce optimism on the derivatives and achieve optimal regret in an adversarial online learning problem. In contrast, existing works either develop optimistic algorithms on the reward (or loss) function as the original UCB strategy (Lai and Robbins,, 1985), or instead use unbiased stochastic gradients and conduct first-order methods for online optimization (Hazan,, 2016).

### v2

@@ 4 Algorithm Design
In this section, we present our core algorithm,C20CB (see Algorithm˜1), which stands for a Closest-To-Zero Confidence Bound strategy that proposes asymptotically optimal prices over differentiated inventory levels and censoring effects.
Algorithm 1 C20CB: Closest-To-Zero Confidence Bound (Main Algorithm)
1:Input: Parameters $a_{\max},b_{\min},b_{\max},p_{\max},c$ and self-derived quantities $\tau,\gamma_{0},C_{a},C_{b},C_{F},C_{G},C_{N},C_{\tau}$.
2://STAGE 1: Pure Exploration for $\tau$ times.
3: Estimate $(\hat{b},\hat{a})\leftarrow\mathit{PureExp}(\tau,p_{\max},\gamma_{0})$ according to Algorithm˜2.
4://STAGE 2: Optimistic Acting
@@
5: Define $\Delta:=(C_{a}+C_{b}\cdot p_{\max})\cdot\frac{1}{\sqrt{\tau}},M:=\lfloor\frac{c}{2\Delta}\rfloor$ and $w_{k}:=2k\Delta,k=-M,-M+1,\ldots,M$.
6: Initialize $(F_{k},N_{k},G_{k},\Delta_{k})\leftarrow\mathit{ConfBoundInit}(\tau,M,\hat{a},\hat{b},c,b_{\max},p_{\max},C_{b},C_{F},C_{G})$ for $k=-M,\ldots,M$, according to Algorithm˜3.
7:for $t=1,2,\ldots,T-\tau-(2M+1)$ do
8:  if $\gamma_{t}\geq\frac{\hat{a}+C_{a}\cdot\frac{1}{\sqrt{\tau}}}{2}+c$ then
9:   Propose $p_{t}\leftarrow\frac{\hat{a}}{2\hat{b}}$ and continue to $t+1$ (without recording feedback).
10:  else
11:   Get $(p_{t},k_{t})\leftarrow\mathit{OptPrice}(M,\gamma_{t},\{(F_{k},N_{k},G_{k},\Delta_{k})\}_{k=-M}^{M},c,\gamma_{0})$ according to Algorithm˜4.
12:   Propose $p_{t}$ as the price, and observe $D_{t}$ and $\mathds{1}_{t}:=\mathds{1}[D_{t}\leq\gamma_{t}]$.
13:   if $k_{t}>$-$\infty$ then
14:    Update
@@ 4.1 Algorithm Design Overview
    STAGE 1: Exploration: During the first $\tau=\sqrt{T}$ rounds, the seller (we) proposes uniformly random prices in the range of $0,p_{\max}]$. By the end of STAGE 1, we obtain $\hat{a}$ and $\hat{b}$ as plug-in estimators of $a$ and $b$ in the following stage.
@@
    STAGE 2: Optimistic Decision: We estimate the derivatives of the revenue function at discretized prices $\{p_{k,t}\}$’s. For each $p_{k,t}$, we not only estimate $r^{\prime}_{t}(p_{k,t})$ but also maintain an error bar of that estimate. At each time $t$, we propose the price whose corresponding error bar covers 0 or closest to 0 if no covering exists.
Algorithm˜1 exhibits several advantageous properties. It is suitable for processing streaming data as the constructions of $\hat{a},\hat{b},\hat{r}^{\prime}_{t}(\cdot)$ are updated incrementally with each new observation (including $e_{i,t},D_{t},\mathds{1}_{t}$) without the need of revisiting any historical data. Additionally, it consumes $\tilde{O}(T^{\frac{5}{4}})$ time complexity and $O(T^{\frac{1}{4}})$ extra space, which are plausible for large $T$. A potential risk of computation might arise on the calculation of $\hat{b}$, where $\sum_{t=1}^{\tau}e_{1,t}-e_{2,t}$ can be 0 with a small but nonzero probability. Although this event does not undermine the high-probability regret guarantee, it might still be harmful to the computational system for numerical experiments. To mitigate this incident in practice, we may either extend STAGE 1 until one non-zero $e_{1,t}-e_{2,t}=1$ is observed, or restart STAGE 1 at $t=\tau$.
Algorithm 2 PureExp: Pure Exploration
1:for $t=1,2,\ldots,\tau$ do
2:  Sample and propose a price $p_{t}\sim U[0,p_{\max}]$ uniformly at random.
3:  Observe demand $D_{t}$, and indicators $e_{i,t}$ as defined in Eq.˜7 for $i=1,2,3$.
4:end for
5: Estimate
$\hat{b}$= $\frac{1}{4p_{\max}\cdot\frac{1}{\tau}\sum_{t=1}^{\tau}\frac{e_{1,t}-e_{2,t}}{\gamma_{t}-\gamma_{0}}},\quad\hat{a}=\frac{1}{\tau}\sum_{t=1}^{\tau}(\hat{b}p_{\max}e_{3,t}+\frac{3\gamma_{t}+\gamma_{0}}{4}).$ (5)
Algorithm 3 ConfBoundInit: Confidence Bound Initialization
1:for $t=1,2,\ldots,2M+1$ do
2:  Let $k_{t}:=-M-1+t$ and propose $p_{t}=\frac{w_{k_{t}}-(\gamma_{t}-\hat{a})}{\hat{b}}$.
3:  Observe $D_{t}$ and $\mathds{1}_{t}:=\mathds{1}[D_{t}<\gamma_{t}]$.
4:  Initialize $F_{k_{t}}\leftarrow\mathds{1}_{t},N_{k_{t}}\leftarrow 1,G_{k_{t}}\leftarrow D_{t}-\gamma_{t}+c,\Delta_{k_{t}}\leftarrow C_{F}\cdot b_{\max}p_{\max}+C_{G}+C_{b}\cdot\frac{1}{\sqrt{\tau}}$.
5:end for
Algorithm 4 OptPrice: Select Optimal Price
1: Initialize $k_{t}\leftarrow M$ as the index of arm to inspect, $\rho_{t}\leftarrow+\infty$ as the smallest absolute value of the derivative estimates we have observed in Time Period $t$ so far.
2:for $k=M,M-1,\ldots,-M+1,-M$ do
3:  Denote $p_{k,t}:=\frac{w_{k}-(\gamma_{t}-\hat{a})}{\hat{b}}$ and $\hat{r}_{k,t}:=\gamma_{0}-c+G_{k}-\hat{b}\cdot p_{k,t}\cdot F_{k}$.
4:  if $\hat{r}_{k,t}-\Delta_{k}\leq 0\leq\hat{r}_{k,t}+\Delta_{k}$ then
5:   Update $k_{t}\leftarrow k,\rho_{t}\leftarrow 0$, and Break.
6:  end if
7:  Let $\rho_{k,t}:=\min\{|\hat{r}_{k,t}-\Delta_{k}|,|\hat{r}_{k,t}+\Delta_{k}|\}$ as the smallest absolute derivative estimate of arm $k$.
8:  if $\rho_{k,t}<\rho_{t}$ then
9:   Update $\rho_{t}\leftarrow\rho_{k,t}$ and $k_{t}\leftarrow k$.
10:  end if
11:end for
12:if $\hat{r}_{k,t}-\Delta_{k}>0,\forall k=-M,-M+1,\ldots,M-1,M$ then
13:  Output $p_{t}\leftarrow\frac{\hat{a}}{2\hat{b}}$ and $k_{t}\leftarrow-\infty.$
14:else
15:  Output $p_{t}\leftarrow p_{k_{t},t}$ and $k_{t}$.
16:end if
Figure 1: The price C20CB proposes based on confidence bounds of $\hat{r}_{k,t}$: (a) If there exist prices whose error bar contain 0, then we propose the largest price among them. (b) If no error bar contains 0 but there does exist at least one below 0, we propose the price whose corresponding error bar is closest to 0. (c) If all error bars are above 0, we propose $p_{t}=\frac{\hat{a}}{2\hat{b}}$.
@@ 4.2 Pure-Exploration to Estimate Parameters from Biased Observations
As shown in Algorithm˜2, we incorporate a uniform exploration phase to estimate $a$ and $b$, bypassing the obstacle caused by demand censoring. This approach is supported by the following insight: When $Y$ is a uniformly distributed random variable within a closed interval $[L,R]$, and $X$ is another random variable, independent to $Y$ and also distributed within $[L,R]$, we have:
@@
Here the second step uses the Law of Total Expectation. Eq.˜6 indicates that we can derive an unbiased estimator of $\mathbb{E}[X]$ through $\mathds{1}[Y\geq X]$ even in the absence of any direct observation of $X$. Looking back to our algorithm, denote $\gamma_{i,t}:=\frac{i\gamma_{t}+(4-i)\gamma_{0}}{4},i=1,2,3$ as the three quarter points of $\gamma_{t}$ for $t=1,2,\ldots,T$, and define
$e_{i,t}=\mathds{1}[D_{t}\geq\gamma_{i,t}],i=1,2,3,t=1,2,\ldots,T.$ (7)
When $p_{t}\sim U[0,p_{\max}]$, we have
$\mathbb{E}[e_{i,t}]$ $=\mathbb{E}[\mathds{1}[a-bp_{t}+N_{t}\geq\gamma_{i,t}]]=\mathbb{E}[\mathbb{E}[N_{t}\geq\gamma_{i,t}-a+bp_{t}|N_{t}]]=\mathbb{E}[\frac{N_{t}-\gamma_{i,t}+a}{bp_{\max}}]=\frac{a-\gamma_{i,t}}{bp_{\max}}.$ (8)
The last equality comes from $\mathbb{E}[N_{t}]=0$. By deploying different $\gamma_{i,t}$ at $i=1,2,3$, we can estimate $a$ and $b$ through the observations of $e_{i,t}$ according to Algorithm˜2, effectively circumventing the censoring effect. A similar technique has been used by [26] to construct an unbiased estimator of valuation instead of the demand, as we are concerned. However, their application of uniform exploration might be sub-optimal as they adopt an exploration-then-exploitation design. In contrast, our algorithm uses this uniform exploration merely as a trigger of further learning. Our tight regret bound indicates that uniform exploration can still contribute to an optimal algorithm for a broad range of online learning instances.
4.3 Optimistic Strategy to Balance Derivatives Estimates v.s. Loss
With $\hat{a}$ and $\hat{b}$ established, we have an estimate of the underlying linear demand $d_{t}(p)=a-bp$. However, we are still unaware of the noise distribution, which is crucial for the current optimal price, since the inventory level $\gamma_{t}$ partially censors the noise.
In order to balance the learning of noise distribution versus the loss of proposing sub-optimal prices, we apply an optimistic strategy in STAGE 2. Usually, an “optimistic strategy” chooses actions as if the best-case scenario within its current confidence region holds, thereby encouraging exploration of potentially rewarding options while still balancing risk.
STAGE 2 involves the following components:
@@
    We discretize the [c,c][-c,c] domain of noise CDF $F(\cdot)$ and its integration $G(\cdot)$ into small intervals of length $2\Delta$, with $\Delta=O(\frac{1}{T^{1/4}})$. At the center of interval $k$ (which is $2k\Delta$) for each $k=-M,-M+1,\ldots,M$ (with $M:=\lfloor\frac{c}{2\Delta}\rfloor$), we maintain independent estimates of $F$ and $G$, including their expectations and high-confidence error bars.
@@
    At each time $t$, we construct a set of discrete prices $\{p_{k,t}\}_{k=-M}^{M}$ such that the quantity $\gamma_{t}-\hat{a}+\hat{b}p_{k,t}$ matches the center of Interval $k$. Given this, we further construct an estimate of each $r^{\prime}_{t}(p_{k,t})$ with plug-in estimators $\hat{a},\hat{b}$ and discrete estimators of $F$ and $G$ for the specific Interval $k$. The estimate includes its expectation and error bar (see Algorithm˜3).
@@
    Since the optimal price $p_{t}^{*}$ satisfies $r^{\prime}_{t}(p_{t}^{*})=0$, we identify the discrete price $p_{k,t}$ where the derivative estimate is “possibly 0” or “closest to 0”. To make this, we design Algorithm˜4 as a component of C20CB and illustrate the process in Figure˜1, which includes the following three cases:
@@
      If there exists some $p_{k,t}$ such that the corresponding error bar (of its derivative estimate) contains 0, we propose the largest price satisfying this condition.
@@
      If there is no $p_{k,t}$ whose corresponding error bar contains 0 (but there exists an error bar below 0), we propose the price whose error bound is closest to 0.
@@
      If all error bars are above 0, indicating that the reward function is monotonically increasing over the “censoring area”, we propose $p_{t}=\frac{\hat{a}}{2\hat{b}}$ to exploit the non-censoring optimal price $\frac{a}{2b}$. In this case, we do not need to record any observations nor to update any parameter/estimate.
@@
Here we provide an intuition of the optimality: On the one hand, the width of the interval can tolerate the error of mapping from $2k\Delta$ to $p_{k,t}$, and the Lipschitzness of $F$ and $G$ ensures that our estimate within each small interval is roughly correct. On the other hand, we can show that the closest-to-zero derivative estimate implies a closest-to-$p_{t}^{*}$ price according to some locally strong convexity in the neighborhood of $p_{t}^{*}$. As we have smoothness on the regret function, we suffer a quadratic loss at $(T\cdot\frac{1}{T^{1/4}})^{2}=O(\sqrt{T})$ cumulatively, which balances the loss of STAGE 1 that costs $O(\tau)=O(\sqrt{T})$ as well. For a rigorous regret analysis, we kindly refer the readers to Section˜5.
Technical Highlights.
This work is the first to introduce optimism on the derivatives and achieve optimal regret in an adversarial online learning problem. In contrast, existing works either develop optimistic algorithms on the reward (or loss) function as the original UCB strategy [37], or instead use unbiased stochastic gradients and conduct first-order methods for online optimization [30].